| 10 - k | = 3
| k - 5 | = 8
what is the value of k that satisfies both equations above ?

Answer by alicealc(293)

| 10 - k | = 3
10 - k = 3 or 10 - k = -3
-k = 3 - 10 or -k = -3 - 10
-k = -7 or -k =-13
k = 7 or k = 13

| k - 5 | = 8
k - 5 = 8 or k - 5 = -8
k = 8 + 5 or k = -8 + 5
k = 13 or k = -3

so, the value of k that satisfies both equation is 13